OpenStudy (littlet):

Is this the correct formula for arc length? Arc length = 2ttr x central angle/360 degrees

OpenStudy (littlet):

@Smartestbruh do you know if it is right?

OpenStudy (phi):

what is 2ttr ?

OpenStudy (littlet):

hold on let me check

OpenStudy (littlet):

I am not sure but it is what i found in my lesson 2TTr

OpenStudy (phi):

ok, that is \[ s = 2 \pi r \frac{\theta}{360} \] the TT is really pi \( \pi\)

OpenStudy (littlet):

Oh ok got it! thank you! idk why i thought it was TT lol

OpenStudy (phi):

people use a Greek letter pi for a "special" number 3.14159.... (it goes on forever)

OpenStudy (littlet):

what is the s= for? is it not arc length?

OpenStudy (littlet):

and oh ok.

OpenStudy (phi):

yes, s is "arc length" In calculus, people use s (short for spatium, Latin for length). Back in the time of Newton, when they invented this stuff, all the educated people used Latin

OpenStudy (littlet):

Oh alright so they are the same thing. ok that makes sense. Well Thank you so much!

OpenStudy (phi):

your equation is correct, as long as you measure the central angle in degrees. If you use radians (another way to measure angles) the arc length formula simplifies to \[ s = r \ \theta\] theta in radians

OpenStudy (littlet):

alright. Thank you.